# Dr.Hair: Reconstructing Scalp-Connected Hair Strands without Pre-Training via Differentiable Rendering of Line Segments

### v1

@@ 1 Introduction
High-quality 3D hair data is essential for depicting realistic human figures in movies, games, and metaverse. However, capturing real hair is notoriously difficult due to its intricate properties, including its elongated shape, overlapping strands, transparency, reflectiveness, and uniformity, which challenge even the most advanced computer vision stereo techniques at a sub-pixel resolution. As a result, hair processing is one of the most formidable tasks in image-based 3D human modeling.
Hair-specific reconstructions methods have been studied for many years [38, 60, 39, 18, 35, 53, 33]. These approaches are characterized by applying a Gabor filter to the hair image to calculate the 2D orientation, which is then combined with 3D measurements for optimization. However, these methods require accurate calibration of the illumination and cameras, making it difficult to scale them to casual shooting environments. In addition, since only the hair’s surface could be measured, it is not easy to estimate the form of the hair connected to the scalp, which is commonly used in industry.
@@ 3.1 Initialization
We leverage gradient domain processing [40, 20] to estimate spatially smooth internal hair flow through volumetric representaiton [66, 23]. A visualizaion of this process is given in Figure 4.
@@
Figure 4: Gradient-domain strand initialization: (a) Boundary condition. Colored region represents $\mathbb{H}$ and $\mathbb{S}$, and black region represents $\mathbb{U}$. (b) Optimized volume sliced by certain planes. Interior is smoothly filled. (c) Extracted strands from the optimized volume.
@@ 3.2 Hierarchical Strand Optimization
A two-stage structure of guide and child hair is employed to express hairstyles. Guide hair grows from each vertex of the scalp mesh, while child hair grows from sampled positions by Sobol sequence [52]. Child hair shape is linearly interpolated with the nearest four guide hairs. 653 guide strands and 50,000 child strands used in this paper. For children, the geometry $\mathscr{G}_{c}=\{\mathscr{V}_{c},\mathscr{F}_{c}\}$ are defined in the same way, and the same rasterization is applied.
@@
We efficiently perform strand optimization following the Coarse-to-Fine strategy. First, the vertex positions of guide hair $\mathscr{V}$ are optimized. Reparameterization with $k=K_{g}$ is applied for guid hair. We rasterize the strands at a thickness close to the actual hair, 0.2 mm. $L_{g}$ is minimized using the Adam optimizer [22] in $I_{g}$ iterations.
@@ 4.1 Synthetic data
Quantitative comparison with existing methods and ablation study are shown in Table 2. Our full pipeline shows better values than the other methods in all criteria. In particular, the much higher recall values indicate that our method successfully recovers internal hair directions that were difficult to handle with existing methods. Effectiveness of each component in the proposed pipeline is also validated. For straight hair with simple internal flow, our strand initialization worked well and showed good scores even without DR. For complex curly hair, the influence of other modules, global optimization, reparameterization, and guide-child hierarchy becomes more pronounced.
@@ 4.2 Real data
Table 2: Quantitative comparison with existing methods and ablation study on synthetic data. P, R, and F1 denote precision, recall, and F1 score, respectively. Higher is better. The lower rows describe the values of our full pipeline and ours without individual modules. w/o DR: DR optimization is not applied, and the intialized strands are evaluated. w/o guide opt.: Child strands are optimized from the beginning of the DR step. w/o reparam.: Reparameterization is disabled. w/o $\mathscr{N}$: Only $\mathscr{N}$ is abandoned in the reparameterization. w/o strand init.: Strands are initialized by straight lines parallel to the normal of the scalp. w/o global opt.: Only gravity heuristic is applied to the initial 3D orientation, and 180° ambiguity is accepted on the other steps.
@@
Figure 10: Left and middle: Re-rendering comparison. Camera image is put on the top for hairstyle reference. The images are ray-traced by Blender Cycles, using the common hair material and lighting. Incorrect orientation of NeuralHaircut made artifacts at meeting points of opposite hair flow, which ours does not have. Right: Volumetric slice of the middle view with certain near and far planes. NeuralHaircut’s hair is concentrated on the surface and scanty inside, whereas our method fills inside uniformly.
@@ 4.3 Limitations
Our method is affected by the input raw mesh quality. If the raw mesh shape has a notable difference from the subject, our strand reconstruction is deteriorated as shown in Figure 12. In addition, the assumption of a smooth flow makes it challenging to handle discontinuous hairstyles such as braids. Moreover, protruding strands can disrupt the overall flow, limiting the accuracy of ours, particularly for highly curly or spiky hair. Finally, as shown in Figure 10, shaded hair color deviates from the actual one because our method does not recover material and lighting.
@@ 6.2 Real data
H3DS is a real-world multi-view dataset for head reconstruction. GT Head meshes was scanned by laser scanner, and independently captured multi-view images surrounding the subject in 360° were registrated against the mesh. About 70 images are provided per subject. Top views are not well captured in both the mesh and the images. Moreover, some views are affected by strong flash lighting that deteriorates image quality. For NeuralHaircut, by following the official implementation setting, 32 clean views manually annotated by the dataset authors were used. For the other methods, all views were inputted.
@@ 6.3 Existing methods’ settings
Table 3: Quantitative comparison with existing methods and ablation study on synthetic data tolerating 180° ambiguity, which have been used in the previous studies. P, R, and F1 denote precision, recall, and F1 score, respectively. Higher is better. The lower rows describe the values of our full pipeline and ours without individual modules. w/o DR: DR optimization is not applied, and the intialized strands are evaluated. w/o guide opt.: Child strands are optimized from the beginning of the DR step. w/o reparam.: Reparameterization is disabled. w/o $\mathscr{N}$: Only $\mathscr{N}$ is abandoned in the reparameterization. w/o strand init.: Strands are initialized by straight lines parallel to the normal of the scalp. w/o global opt.: Only gravity heuristic is applied to the initial 3D orientation, and 180° ambiguity is accepted on the other steps.
@@ 7.1 Additional comparison and ablations on synthetic data
Figure 16: Additional results on studio data 1 / 2. From left to right, top view image, top view guide with 3D orientation, top view child with Blender Cycles shading and 3D orientation, side view image, side view guide with random color, and side view child with Blender Cycles shading and random color are shown. From top to bottom, the subjects with, short hair, long hair, half bald head, and short-tied hair are displayed. Our method enables realistic reconstruction for a wide range of hairstyles in the wild.
@@
Figure 18: Comparison of physics simulation with head motion. Full sequences are available in the supplemental video. Starting from the reconstructed strands, gravity, hair stiffness and head rotation are applied to NeuralHaircut and ours. On top two rows, the subjects of studio data in Figure 1 and 10 of the main paper are shown. The bottom row shows the subject of H3DS in the top row of Figure 7 of the main paper. Thanks to correct hair growing direction, our hair shows more natural behavior under strong head movement. The original scalp shapes and hair root positions are kept while the head model is replaced for privacy protection.
@@
Figure 23: Comparison of scalp normal and our $S(p_{s})$. (a) Normals at scalp, $n_{s}(p_{s})$. Growing directions at the side and back are different from real humans. (b) Our hair orientation at scalp $S(p_{s})$. Strands can grow more naturally at the side and back.
@@ 7.2 Additional results on studio data
7.3 Addtional physics simulation results
In Figure 18, the physics simulations with head movement were compared in time series. The full sequence is available in the supplementary video. The natural behavior of our strands indicate that our method is capable of reconstructing simulation-ready hair strands.
@@ 8.1 Comparison with existing AAs
In the validation, the width of our strand is set to 0.2 mm, which is often thinner than one pixel. So, in this validation, we tested root thickness in 1.0, 0.8, and 0.6 pixels. Thin width is prone to cause jumping pixels by the nature of rasterization.
@@
Although the proposed pipeline utilizes the AA for fine-tuning following initialization, it possesses sufficient capability for growth. The depiction of the “CVPR” drawing with strands is presented in Figure 22. Each GT letter is constructed using one, two, two, and three bundles for “C”, “V”, “P”, and “R” respectively, wherein a bundle comprises 500 strands with 50 segments functioning as child hair, each capable of independent movement. Target images were rendered from random views of GT hairs on a per-letter basis. GT hairs were shortened to 10% of their original length as starting values for optimization utilizing DR. Each letter was optimized independently by Adam optimzier. The loss comprised $L_{m}$, $L_{o}$, $R_{root}$, $R_{c}$, alongside regularizer for equalizing segment lengths. For reparameterization, $k$NN with $\mathscr{N}=10$ was performed at 50% of the original length to address artifacts occurring at junctions. The “CVPR” drawing demonstrates that our AA can optimize complex shapes effectively.
@@ 9.4 Hair mask for DR loss
9.5 Module-level peformance measurement

### v2

@@ 1 Introduction
High-quality 3D hair data is essential for depicting realistic human figures in movies, games, and metaverse. However, capturing real hair is notoriously difficult due to its intricate properties, including its elongated shape, overlapping strands, transparency, reflectivity, and uniformity, which challenge even the most advanced computer vision stereo techniques at a sub-pixel resolution. As a result, hair processing is one of the most formidable tasks in image-based 3D human modeling.
Hair-specific reconstructions methods have been studied for many years [38, 60, 39, 18, 35, 53, 33]. These approaches are characterized by applying a Gabor filter to the hair image to calculate the 2D orientation, which is then combined with 3D measurements for optimization. However, these methods require accurate calibration of the illumination and cameras, making it difficult to scale them to casual shooting environments. In addition, since only the hair’s surface can be measured, it is not easy to estimate the form of the hair connected to the scalp, which is commonly used in industry.
@@ 3.1 Initialization
We leverage gradient domain processing [40, 20] to estimate spatially smooth internal hair flow through volumetric representaiton [66, 23]. A visualization of this process is given in Figure 4.
@@
Figure 4: Gradient-domain strand initialization: (a) Boundary condition. Colored region represents $\mathbb{H}$ and $\mathbb{S}$, and black region represents $\mathbb{U}$. (b) Optimized volume sliced by certain planes. The Interior is smoothly filled. (c) Extracted strands from the optimized volume.
@@ 3.2 Hierarchical Strand Optimization
A two-stage structure of guide and child hair is employed to express hairstyles. Guide hair grows from each vertex of the scalp mesh, while child hair grows from sampled positions by Sobol sequence [52]. Child hair shape is linearly interpolated with the nearest four guide hairs. 653 guide strands and 50,000 child strands are used in this paper. For children, the geometry $\mathscr{G}_{c}=\{\mathscr{V}_{c},\mathscr{F}_{c}\}$ are defined in the same way, and the same rasterization is applied.
@@
We efficiently perform strand optimization following the Coarse-to-Fine strategy. First, the vertex positions of guide hair $\mathscr{V}$ are optimized. Reparameterization with $k=K_{g}$ is applied for guide hair. We rasterize the strands at a thickness close to the actual hair, 0.2 mm. $L_{g}$ is minimized using the Adam optimizer [22] in $I_{g}$ iterations.
@@ 4.1 Synthetic data
Quantitative comparison with existing methods and ablation study are shown in Table 2. Our full pipeline shows better values than the other methods in all criteria. In particular, the much higher recall values indicate that our method successfully recovers internal hair directions that were difficult to handle with existing methods. The effectiveness of each component in the proposed pipeline is also validated. For straight hair with simple internal flow, our strand initialization worked well and showed good scores even without DR. For complex curly hair, the influence of other modules, global optimization, reparameterization, and guide-child hierarchy becomes more pronounced.
@@ 4.2 Real data
Table 2: Quantitative comparison with existing methods and ablation study on synthetic data. P, R, and F1 denote precision, recall, and F1 score, respectively. Higher is better. The lower rows describe the values of our full pipeline and ours without individual modules. w/o DR: DR optimization is not applied, and the initialized strands are evaluated. w/o guide opt.: Child strands are optimized from the beginning of the DR step. w/o reparam.: Reparameterization is disabled. w/o $\mathscr{N}$: Only $\mathscr{N}$ is abandoned in the reparameterization. w/o strand init.: Strands are initialized by straight lines parallel to the normal of the scalp. w/o global opt.: Only gravity heuristic is applied to the initial 3D orientation, and 180° ambiguity is accepted on the other steps.
@@
Figure 10: Left and middle: Re-rendering comparison. The camera image is put on the top for hairstyle reference. The images are ray-traced by Blender Cycles, using the common hair material and lighting. Incorrect orientation of NeuralHaircut made artifacts at meeting points of opposite hair flow, which ours does not have. Right: Volumetric slice of the middle view with certain near and far planes. NeuralHaircut’s hair is concentrated on the surface and scanty inside, whereas our method fills inside uniformly.
@@ 4.3 Limitations
Our method is affected by the input raw mesh quality. If the raw mesh shape has a notable difference from the subject, our strand reconstruction deteriorates, as shown in Figure 12. In addition, the assumption of a smooth flow makes it challenging to handle discontinuous hairstyles such as braids. Moreover, protruding strands can disrupt the overall flow, limiting our accuracy, particularly for highly curly or spiky hair. Finally, as shown in Figure 10, shaded hair color deviates from the actual one because our method does not recover material and lighting.
@@ 6.2 Real data
H3DS is a real-world multi-view dataset for head reconstruction. GT Head meshes were scanned by a laser scanner, and independently captured multi-view images surrounding the subject in 360° were registrated against the mesh. About 70 images are provided per subject. Top views are not well captured in both the mesh and the images. Moreover, some views are affected by strong flash lighting that deteriorates image quality. For NeuralHaircut, by following the official implementation setting, 32 clean views manually annotated by the dataset authors were used. For the other methods, all views were inputted.
@@ 6.3 Existing methods’ settings
Table 3: Quantitative comparison with existing methods and ablation study on synthetic data tolerating 180° ambiguity, which have been used in the previous studies. P, R, and F1 denote precision, recall, and F1 score, respectively. Higher is better. The lower rows describe the values of our full pipeline and ours without individual modules. w/o DR: DR optimization is not applied, and the initialized strands are evaluated. w/o guide opt.: Child strands are optimized from the beginning of the DR step. w/o reparam.: Reparameterization is disabled. w/o $\mathscr{N}$: Only $\mathscr{N}$ is abandoned in the reparameterization. w/o strand init.: Strands are initialized by straight lines parallel to the normal of the scalp. w/o global opt.: Only gravity heuristic is applied to the initial 3D orientation, and 180° ambiguity is accepted on the other steps.
@@ 7.1 Additional comparison and ablations on synthetic data
Figure 16: Additional results on studio data 1 / 2. From left to right, top view image, top view guide with 3D orientation, top view child with Blender Cycles shading and 3D orientation, side view image, side view guide with random color, and side view child with Blender Cycles shading and random color are shown. From top to bottom, the subjects with short hair, long hair, half bald head, and short-tied hair are displayed. Our method enables realistic reconstruction for a wide range of hairstyles in the wild.
@@
Figure 18: Comparison of physics simulation with head motion. Full sequences are available in the supplemental video. Starting from the reconstructed strands, gravity, hair stiffness and head rotation are applied to NeuralHaircut and ours. On the top two rows, the subjects of studio data in Figures 1 and 10 of the main paper are shown. The bottom row shows the subject of H3DS in the top row of Figure 7 of the main paper. Thanks to the correct hair growing direction, our hair shows more natural behavior under strong head movement. The original scalp shapes and hair root positions are kept while the head model is replaced for privacy protection.
@@
Figure 23: Comparison of scalp normal and our $S(p_{s})$. (a) Normals at the scalp, $n_{s}(p_{s})$. Growing directions at the side and back are different from real humans. (b) Our hair orientation at scalp $S(p_{s})$. Strands can grow more naturally at the side and back.
@@ 7.2 Additional results on studio data
7.3 Additional physics simulation results
In Figure 18, the physics simulations with head movement were compared in time series. The full sequence is available in the supplementary video. The natural behavior of our strands indicates that our method is capable of reconstructing simulation-ready hair strands.
@@ 8.1 Comparison with existing AAs
In the experiments of the main paper, the width of our strand is set to 0.2 mm, which is often thinner than one pixel. So, in this validation, we tested root thickness in 1.0, 0.8, and 0.6 pixels. Thin width is prone to cause jumping pixels by the nature of rasterization.
@@
Although the proposed pipeline utilizes the AA for fine-tuning following initialization, it possesses sufficient capability for growth. The depiction of the “CVPR” drawing with strands is presented in Figure 22. Each GT letter is constructed using one, two, two, and three bundles for “C”, “V”, “P”, and “R” respectively, wherein a bundle comprises 500 strands with 50 segments functioning as child hair, each capable of independent movement. Target images were rendered from random views of GT hairs on a per-letter basis. GT hairs were shortened to 10% of their original length as starting values for optimization utilizing DR. Each letter was optimized independently by Adam optimizer. The loss comprised $L_{m}$, $L_{o}$, $R_{root}$, $R_{c}$, alongside regularizer for equalizing segment lengths. For reparameterization, $k$NN with $k=10$ was performed for $\mathscr{N}$ at 50% of the original length to address artifacts occurring at junctions. The “CVPR” drawing demonstrates that our AA can optimize complex shapes effectively.
@@ 9.4 Hair mask for DR loss
9.5 Module-level performance measurement